OpenStudy (anonymous):

A recursive function is shown below: f(1) = 4 and f(n) = f(n - 1) - 8; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? a. 4, -4, -12, -20, -28, ... b. 4, 4, 12, 20, 28, ... c. 4, 12, 20, 28, 36, ... d. 4, -12, -20, -28, -36, ...

OpenStudy (perl):

did you try to plug in

OpenStudy (anonymous):

1 w4snt sur3 wh3r3 to plug th3m 1n.

OpenStudy (perl):

f(1) = 4 then f(2) = f(2-1) - 8 = f(1) - 8 , now substitute

OpenStudy (anonymous):

f(2)= -4 ?

OpenStudy (perl):

correct

OpenStudy (perl):

now use f(2) to find f(3)

OpenStudy (anonymous):

th4nks!

OpenStudy (perl):

f(3) = f(2) - 8

OpenStudy (anonymous):

-12